OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (tyteen4a03):

I've already shown you how to solve these type of questions earlier. Try it yourself :)

OpenStudy (anonymous):

I prefer to solve by addition/elimination... I am going to eliminate x so we can find y first. 3x - 6y = 12 --->(2)3x - 6y = 12 -2x + 3y = 6 --> (3)-2x + 3y = 6 ----------------------------- 6x - 12y = 24 (result from multiplying by 2) -6x + 9y = 18 (result from multiplying by 3) ----------------------------- 0 - 3y = 42 -3y = 42 -3y/-3 = 42/-3 y = - 14